OpenStudy (anonymous):

On a number line, find the midpoint of a segment with the following endpoints: A = -10 and B = 20 1) 10 2) 5 3) 12 4) 5.5

OpenStudy (amistre64):

20-(-10) = 30 for a total distance 30/2 = 15 equals distance to the midpoint -10 + 15 = 5 id go with option 2